OpenStudy (anonymous):

Mr. Brownwood invests a certain amount of money at 9% interest and $1,800 more than that amount in another account at 11% interest. At the end of one year, he earned a total of $818 in interest. How much money was invested in each account? a. $3,500 at 9%; $4,300 at 11% b. $3,400 at 9%; $3,200 at 11% c. $3,100 at 9%; $4,900 at 11%

OpenStudy (btaylor):

0.09x + 0.11(x+1800)=818 .09x + .11x + 198 = 818 0.20x=620 x = 3100 x+1800 = 4900 so the answer is (C)

OpenStudy (anonymous):

im not sure if i get it...

OpenStudy (btaylor):

you multiply each interest rate by the amount you invested at that rate, then add them together to get the total interest earned.

OpenStudy (anonymous):

i guess i kinda get it... >.<

OpenStudy (anonymous):

anyway thanks